OpenStudy (anonymous):

What is (8.7)/(8.1)?

OpenStudy (unklerhaukus):

\[\frac{8.7}{8.1}=\frac{8.1+0.6}{8.1}=\frac{8.1}{8.1}+\frac {0.6}{8.1}\]

OpenStudy (unklerhaukus):

times the top and bottom of the fraction by 10 to remove the decimal point